Suppose an individual had a utility function given by: U=X^4*Y^1.

Calculate this individual's Marginal Rate of Substitution (MRSxy) when they have a bundle with 5 units of Good X and 0.25 units of Good Y.

(Round to the nearest decimal place if necessary.)
Answer #1

Utility function is given by U = X^4 Y^1

MRSxy = MUx/MUy

= dU/dx divided by dU/dy

= 4X^3 Y^1 / X^4 * 1 * Y^0

= 4X^3 Y / X^4

= 4Y/X

At (X = 5, Y = 0.25), MRS = 4*0.25 / 5 = 0.20

Hence MRS is 0.20.